OpenStudy (anonymous):

Generate the first 5 terms of this sequence:

OpenStudy (anonymous):

f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 0, -1, 1, 0, 2 0, 1, 1, 2, 3 0, 1, 2, 2, 3 0, 1, 1, 2, 2

OpenStudy (anonymous):

to find f(3) just add f(1) and f(2), f(3)=0+1=1, for f(4) add f(2) and f(3), f(4)=1+1=2, then for the last add f(3) and f(4), f(5)=1+2=3 so it is 0,1,1,2,3